what is the slope of 5x+7
Answer by sdmmadam@yahoo.com(530)

what is the slope of 5x+7?
The question is not very clear.
1)If the question is what is the slope of y =5x +7
Here is the answer and the formula to remember
The given equation is y = 5x + 7
then the slope of the straight line is 5.
Note: An equation of the form
y = mx + c ----(*)
is called the equation to a line in slope and y-intercept form.
When the coefficient of y is (1) on one side say LHS of the equation, the coefficient of x on the other side (that is RHS)gives the slope of the straight line. And the free constant on the RHS is the y-intercept of the straightline.
Note: Given the general form lx +my +n =0 of a straight line how to find the slope and y-intercept of the straight line.
Consider ux+ vy + n = o ----(1)
Retain the term containing y that is (vy) here on the LHS and transfer the other terms to the other side(changing sign while changing side)
We get vy = -ux -n
Dividing by v through out
y = (-u/v)x +(-n/v)
and comparing it with the standard slope and y-intercept form,
the slope m = (-u/v) and y-intercept = (-n/v)
Note: What happens if the coefficient v is zero?
If v = o then (vy) = 0 (as zero multiplied by anything is zero)
and the equation contains only the x-term and the constant term which means the equation will be of the form x = constant
How?
If v=0, then the given line (1) becomes ux + n= 0 which implies ux = -n and in that case for u not equal to zero, x = (-n/u)= a constant and x=constant is a line parallel to the y-axis and has slope = infinity (as the y -axis as well as any line parallel to the y-axis makes an angle 90 degrees with the x-axis and tangent of (90 degrees) is infinity(the reciprocal of zero)
2) If the given equation is 5x +7 = 0, then you can understand from the foot notes I have given above that the given line is a line parallel to the x axis
and the slope is infinity.
Note: The slope or the gradient of a straight line is defined as the tangent of the angle which the line makes with the x-axis, the angle being measured in the anti-clockwise direction.